QUESTION 1

  1. Determine Tyrion Company’s total variable cost (TVC).

Sales (125,000 units)

$4,925,000

Total fixed operating costs

$2,687,500

Total fixed financing costs

$800,000

Operating breakeven quantity of sales

75,000 units

a.

$256,667

b.

$376,667

c.

$445,833
Answer #1

If it is helpful, please rate the answer and if any doubt arises let me know

Correct answer is: c. $445833
Workings:
Let Variable cost per unit be X
Sales Price per unit ($4925000 / 125000) = $                                     39
Variable cost per unit = X
Contribution margin per unit = $39 - X
Total Fixed operating Cost = $                       26,87,500
Operating Break - even quantity of sales = Total Fixed operating Cost / Contribution margin per unit
75000 = $2687500 / ($39 - X)
$39 - X = $2687500 / 75000
$39 - X = $                               35.83
X = $                                 3.57
Total Variable cost (TVC) ($3.57 X 125000) = $                         4,45,833